(2) A 3.8g metal made from a gun residue containing iron is dissolved in 100cm acid solution. The solution formed contains (F can you please explain why you do esch step too.
Answer #1

The moles of MnO4- used, which are the same as Fe + 3, are calculated:

n MnO4- = M * V = 0.02 M * 0.0113 L = 0.000226 mol

The mass of Fe in 15 cm3 is calculated:

m Fe = n * MM = 0.000226 * 55.8 = 0.013 g

The mass in the total diluted volume is calculated:

m1 = 0.013 * 100/15 = 0.087 g

The percentage is calculated:

% Fe = 0.087 * 100 / 3.8 = 2.29%

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.